# Rajamanickam

### v5

@@ Abstract
Strong anchoring boundary conditions are conventionally modelled by imposing Dirichlet conditions on the order parameter in Landau–de Gennes theory, neglecting the finite surface energy of realistic anchoring. This work revisits the strong anchoring limit for nematic liquid crystals in confined two-dimensional domains. By explicitly retaining a Rapini–Papoular surface energy and adopting a scaling where the extrapolation length $l_{ex}$ is comparable to the coherence length $\xi$, we analyse both the small-domain ($\varepsilon=h/\xi\to 0$; $h$ is the domain size) and Oseen–Frank $(\varepsilon\to\infty$) asymptotic regimes. In the small-domain limit, the leading-order equilibrium solution is given by the average of the boundary data, which can vanish in symmetrically frustrated geometries, leading to isotropic melting. In the large-domain limit, matched asymptotic expansions reveal that surface anchoring introduces an $O(1/\varepsilon)$ correction to the director field near boundaries, in contrast to the $O(1/\varepsilon^{2})$ correction predicted by Dirichlet conditions. The analysis captures the detailed structure of interior and boundary defects, showing that mixed (Robin-type) boundary conditions yield smoother defect cores and more physical predictions than rigid Dirichlet conditions. Importantly, a stability analysis based on the intrinsic-basis decomposition shows that even small deviations of $O(1/\varepsilon^{2})$ from perfect scalar order can destabilise some Oseen–Frank solutions, indicating that only a restricted subset of solutions are realised as stable equilibria in the full Landau–de Gennes framework. Numerical solutions for square and circular wells with tangential anchoring illustrate the differences between the two boundary condition treatments, particularly in defect morphology. These results demonstrate that a consistent treatment of anchoring energetics, together with stability considerations, is essential for accurate modelling of nematic equilibria in micro- and nano-scale confined geometries.
@@ 2 Remarks on the strong anchoring condition in three dimensions
To place these regimes in context, we note that in practice $l_{ex}\sim 10^{-8}-10^{-5}$[ravnik2009landau], with the strongest achievable anchoring corresponding to a few tens of nanometers. By contrast, the coherence length $\xi\sim 10^{-8}m$ [de1993physics] is typically only a few ångströms and rarely exceeds tens of nanometers even very close to the nematic–isotropic transition [^1] Strictly speaking, very close to the nematic–isotropic transition, fluctuation effects eventually invalidate the mean-field Landau–de Gennes description.. For example, using typical material parameters, $L=15pN$, $B=3.7\times 10^{5}Nm^{-2}$ and $C=2.4\times 10^{5}Nm^{-2}$ [rajamanickam2025colloidal], one finds $\xi=2.7\times 10^{-8}m$ at the transition point. Consequently, the regime $l_{ex}\ll\xi$ is seldom realised in practice. At this stage, it is worth noting that most studies in the literature impose Dirichlet conditions even within localized boundary defects corresponding to this infeasible regime.
@@ 3 Problem formulation in the reduced two-dimensional Landau–de Gennes framework
with $\boldsymbol{\nu}$ being the unit outward normal to $\partial\Omega$. The formulation for the stability analysis is provided in Appendi x A.
@@ 4 Small-domain limit, $\varepsilon\to 0$
$q_{1}^{(0)}=\frac{1}{|\partial\Omega|}\oint_{\partial\Omega}q_{1,b}\,dl,\qquad q_{2}^{(0)}=\frac{1}{|\partial\Omega|}\oint_{\partial\Omega}q_{2,b}\,dl.$ (18)
@@ 5 Oseen–Frank limit, $\varepsilon\to\infty$
Equilateral triangle [han2020reduced, rajamanickam2026nematic] 44 three +1+1-vertex defects & one $\tfrac{1}{2}-$interior defect
@@
$\frac{\partial s}{\partial\nu}=\varepsilon\gamma[s-\cos(2\varphi-2\varphi_{b})],\qquad s\frac{\partial\varphi}{\partial\nu}=\varepsilon\gamma\sin(2\varphi-2\varphi_{b})\quad\text{on}\quad\partial\Omega.$ (33)
@@ 5.1 Outer solution
$s_{1}=0,\quad\varphi_{1}=\frac{1}{2\gamma}\frac{\partial\varphi_{0}}{\partial\nu}\quad\text{on}\quad\partial\Omega.$ (38)
@@
$s_{2}=-\frac{1}{2\gamma^{2}}\left(\frac{\partial\varphi_{0}}{\partial\nu}\right)^{2},\quad\varphi_{2}=\frac{1}{2\gamma}\frac{\partial\varphi_{1}}{\partial\nu}\quad\text{on}\quad\partial\Omega.$ (40)
@@
The first-order correction exists only because $\gamma\sim O(1)$. In studies where surface energy is neglected $(\gamma\to\infty)$, $\varphi_{1}=0$ and the first correction vanishes. Properly accounting for surface energy introduces an $O(1/\varepsilon)$ correction to the director field, in contrast to the $O(1/\varepsilon^{2})$ prediction from purely Dirichlet boundary conditions [nguyen2013refined, di2020landau]. The scalar order parameter, however, satisfies
$s\approx 1-2\varepsilon^{-2}|\nabla\varphi_{0}|^{2}+\cdots$ (41)
The mismatch between bulk $s_{2}$ and the boundary $s_{2}$ arises from $4\gamma^{2}\neq 1$ and ${\partial\varphi_{0}}/{\partial\tau}\neq 0$, indicating a weak boundary layer in $s$. The director field $\mathbf{n}=(\cos\varphi,\sin\varphi)^{T}$ admits the expansion
@@ 5.3 Structure of a boundary defect
$\frac{S_{0}}{\gamma\eta}\frac{\partial\Phi_{0}}{\partial\theta}=\sin(2\Phi_{0}-2\varphi_{b})\quad\text{on}\quad\theta=\pm\alpha,$ (67)
@@
5.4 Remarks on the stability
Although the reduced two-dimensional Landau–de Gennes framework formally contains the Oseen–Frank theory as the limiting case $s_{0}\to 1$, this limit is structurally delicate from the perspective of stability. As shown in Appendix A, linear perturbations about a leading-order Oseen–Frank equilibrium naturally decompose into modes associated with variations in the magnitude of $\mathbf{Q}$ and rotations of the director field. In the strict Oseen–Frank limit, director perturbations are elastically stabilised, and the corresponding equilibria are always local minima with respect to director variations.
However, even a small departure from perfect order, $1-s\sim O(\varepsilon^{2})$, removes the bulk restoring force for director rotations and permits coupling to magnitude fluctuations. This is precisely captured by the higher-order corrections calculated from the Euler–Lagrange equations, where $s=1-2\varepsilon^{-2}|\nabla\varphi_{0}|^{2}+\cdots$ and $\mathbf{n}=\mathbf{n}_{0}+\varepsilon^{-1}\varphi_{1}\mathbf{t}_{0}+\varepsilon^{-2}(\varphi_{2}\mathbf{t}_{0}-\tfrac{1}{2}\varphi_{1}^{2}\mathbf{n}_{0})+\cdots$. Consequently, the higher-order terms are not merely refinements; they determine which Oseen–Frank solutions survive as stable equilibria of the full Landau–de Gennes energy and highlight the special sensitivity of defect cores and regions of reduced order. The intrinsic-basis stability analysis in Appendix A makes this restriction precise and demonstrates how coupling between magnitude and director fluctuations may govern the selection of admissible solutions. This does not imply that all deviations from $s=1$ lead to instability; rather, it implies that only a restricted class of Oseen–Frank solutions can persist as stable equilibria of the full Landau–de Gennes energy.
@@ Acknowledgment
The Eigenvalue problem in a natural intrinsic basis
To analyse stability more transparently, it is convenient to project perturbations onto a local intrinsic frame aligned with the unperturbed $\mathbf{Q}$-tensor. The matrix $\bar{\mathbf{q}}\otimes\bar{\mathbf{q}}$ is of rank one, with eigenvalues $(\frac{1}{4},0)$ and corresponding normalised eigenvectors
$\mathbf{e}_{\parallel}=\frac{\bar{\mathbf{q}}}{|\bar{\mathbf{q}}|},\qquad\mathbf{e}_{\perp}=\frac{1}{|\bar{\mathbf{q}}|}\begin{pmatrix}-\bar{q}_{2}\\ \bar{q}_{1}\end{pmatrix}$. (83)
Any perturbation can then be decomposed as
$\hat{\mathbf{q}}=\hat{q}_{{}_{\parallel}}\mathbf{e}_{\parallel}+\hat{q}_{\perp}\mathbf{e}_{\perp}$ (84)
where $\hat{q}_{\parallel}$ represents variations in the magnitude of the $\mathbf{Q}$-tensor and $\hat{q}_{\perp}$ corresponds to reorientation of the director. The above decomposition decouples the parallel and perpendicular modes in the bulk potential, though the elastic term becomes more intricate due to the spatial variation of the basis vectors.
The local rotation of the unperturbed director is captured by the one-form
$\boldsymbol{\omega}=\mathbf{e}_{\perp}\cdot\nabla\mathbf{e}_{\parallel}=\frac{\bar{q}_{1}\nabla\bar{q}_{2}-\bar{q}_{2}\nabla\bar{q}_{1}}{|\bar{\mathbf{q}}|^{2}}=2\nabla\bar{\varphi}$ (85)
so that $\nabla e_{\parallel}=\boldsymbol{\omega}\mathbf{e}_{\perp}$ and $\nabla e_{\perp}=-\boldsymbol{\omega}\mathbf{e}_{\parallel}$. Projecting the Hessian equation, $\mathbf{H}_{\bar{\mathbf{q}}}\hat{\mathbf{q}}=\sigma\hat{\mathbf{q}}$, onto the intrinsic frame yields
$\Delta_{\boldsymbol{\omega}}\begin{bmatrix}\hat{q}_{\parallel}\\ \hat{q}_{\perp}\end{bmatrix}-\varepsilon^{2}\begin{bmatrix}12|\bar{\mathbf{q}}|^{2}-1&0\\ 0&4|\bar{\mathbf{q}}|^{2}-1\end{bmatrix}\begin{bmatrix}\hat{q}_{\parallel}\\ \hat{q}_{\perp}\end{bmatrix}=-\frac{\sigma}{2}\begin{bmatrix}\hat{q}_{\parallel}\\ \hat{q}_{\perp}\end{bmatrix}$ (86)
where $\Delta_{\boldsymbol{\omega}}$ is the covariant Laplacian, defined by
$\Delta_{\boldsymbol{\omega}}$ $=\mathbf{I}(\nabla^{2}-|\boldsymbol{\omega}|^{2})+\mathbf{J}(2\boldsymbol{\omega}\cdot\nabla+\nabla\cdot\boldsymbol{\omega}),\qquad\mathbf{J}=\begin{pmatrix}0&-1\\ 1&0\end{pmatrix}$, (87)
$=\mathbf{I}(\nabla^{2}-4|\nabla\bar{\varphi}|^{2})+\mathbf{J}(4\nabla\bar{\varphi}\cdot\nabla+2\nabla^{2}\bar{\varphi}).$ (88)
As we can see, the $90^{\circ}$-rotation matrix $\mathbf{J}$ couples the parallel and perpendicular modes. Although $\mathbf{J}$ is skew-symmetric, the operator $\mathbf{J}(2\boldsymbol{\omega}\cdot\nabla+\nabla\cdot\boldsymbol{\omega})$ is self-adjoint, as can be verified directly, ensuring that the eigenvalues remain real. The first term, $\mathbf{I}\nabla^{2}$, represents the standard elastic penalty and acts to smooth perturbations. The second term, $4|\nabla\bar{\varphi}|^{2}\mathbf{I}$, acts like a potential (see below), rendering perturbations more difficult to sustain in regions where the background director varies strongly. The third term, $2\mathbf{J}\boldsymbol{\omega}\cdot\nabla$, captures how perturbations in one mode (say, parallel amplitude changes) get converted to the other mode (perpendicular director angle changes) as one follows the background director rotation. The fourth term, $\mathbf{J}\nabla\cdot\boldsymbol{\omega}$, acts like a local source that converts between perturbation modes based on the local accumulation of director rotation and plays a significant role in defect cores, where $\nabla^{2}\bar{\varphi}\neq 0$ in general. Together, these four contributions fully encode the interplay between director rotation, mode coupling, and elastic smoothing in the linearised dynamics around the equilibrium state.
The bulk potential terms in the Hessian are diagonal in this basis. Since $|\bar{\mathbf{q}}|^{2}\in[0,1/4]$, we have $12|\bar{\mathbf{q}}|^{2}-1\in[-1,2]$ and $4|\bar{\mathbf{q}}|^{2}-1\in[-1,0]$. The parallel term $12|\bar{\mathbf{q}}|^{2}-1$ thus stabilises magnitude fluctuations when $\bar{s}$ is close to unity, but can destabilize in low-order regions such as defect cores. The perpendicular term $4|\bar{\mathbf{q}}|^{2}-1$ governs how director rotations couple to the local degree of order. In perfectly ordered regions, where $\bar{s}=1$, this term is zero, it vanishes, reflecting rotational symmetry and producing soft elastic modes. However, even a slight violation of this condition, $1-\bar{s}\sim O(\varepsilon^{2})$, renders this term negative, so that the bulk potential provides no restoring force for director rotations, which must then rely primarily on elasticity for stability. Together, these effects imply that defect cores are inherently vulnerable to instabilities in both magnitude and orientation, a phenomenon absent in the director‑only Oseen–Frank approximation.
Director stability in the Oseen–Frank limit
The results for the Oseen–Frank limit is simple. For brevity, consider only the leading-order Oseen–Frank problem described in (35)-(36) and drop the subscript ”0”. Since $\bar{s}^{2}=4|\bar{\mathbf{q}}|^{2}=1$ and $\nabla^{2}\bar{\varphi}=0$, the eigenvalue equation (86) simplifies to
$\Delta_{\boldsymbol{\omega}}\begin{bmatrix}\hat{q}_{\parallel}\\ \hat{q}_{\perp}\end{bmatrix}-\varepsilon^{2}\begin{bmatrix}2&0\\ 0&0\end{bmatrix}\begin{bmatrix}\hat{q}_{\parallel}\\ \hat{q}_{\perp}\end{bmatrix}=-\frac{\sigma}{2}\begin{bmatrix}\hat{q}_{\parallel}\\ \hat{q}_{\perp}\end{bmatrix}$ (89)
with $\Delta_{\boldsymbol{\omega}}=\mathbf{I}(\nabla^{2}-4|\nabla\bar{\varphi}|^{2})+4\mathbf{J}\nabla\bar{\varphi}\cdot\nabla$. However, in the Oseen–Frank limit, $\varepsilon\to\infty$, the linear perturbation must remain on the manifold $s=1$ because $\hat{q}_{\parallel}=O(\varepsilon^{-2})$, so that only the perpendicular component is dynamically relevant. The resulting eigenvalue equation reads
$\nabla^{2}\hat{q}_{\perp}+(\tfrac{1}{2}\sigma-4|\nabla\bar{\varphi}|^{2})\hat{q}_{\perp}=0$ (90)
Introducing $\hat{q}_{\perp}=2\hat{\varphi}$, with $\hat{\varphi}$ denoting the perturbation of the director field, one obtains
$\nabla^{2}\hat{\varphi}+(\tfrac{1}{2}\sigma-4|\nabla\bar{\varphi}|^{2})\hat{\varphi}=0,\qquad\hat{\varphi}|_{\partial\Omega}=0.$ (91)
The term $4|\nabla\bar{\varphi}|^{2}$ in the Hessian arises from the intrinsic curvature of the uniaxial constraint manifold and therefore persists in the Oseen–Frank limit, even though it does not appear in the leading-order Euler–Lagrange equations. This eigenvalue problem is formally analogous to the two-dimensional Schrödinger equation with an infinite barrier on $\partial\Omega$, i.e.,
$\nabla^{2}\psi+\frac{2m}{\hbar^{2}}(E-V)\psi=0,\qquad\psi|_{\partial\Omega}=0,$ (92)
under the correspondence
$\hat{\varphi}\longmapsto\psi,\qquad 1\longmapsto\frac{2m}{\hbar^{2}},\qquad\frac{\sigma}{2}\longmapsto E,\qquad 4|\nabla\bar{\varphi}|^{2}\longmapsto V.$ (93)
Since the “potential” $4|\nabla\bar{\varphi}|^{2}\geq 0$, it follows that the “energy” $\sigma/2\geq 0$, indicating that the base state is linearly stable. In fact, the spectrum is discrete and can be ordered as $0<\sigma_{1}<\sigma_{2}\leq\sigma_{3}\leq\cdots$ with $\sigma_{n}\geq\lambda_{n}$, where $\nabla^{2}\phi_{n}=\tfrac{1}{2}\lambda_{n}\phi_{n}$, $\phi_{n}|_{\partial\Omega}=0$. This reflects the strength and limitation of the Oseen–Frank approximation because elastic stability of the director field is captured while instabilities associated with variations in the magnitude of $\mathbf{Q}$ are neglected.
In the full Landau–de Gennes theory, however, variations in the scalar order parameter $s$ can generate negative eigenvalues associated with defect–core instabilities. Within such cores, both $\hat{q}_{\parallel}$ and $\hat{q}_{\perp}$ must be retained, and the reduced Oseen–Frank description provides only an approximate account of the stability landscape. Consequently, Oseen–Frank equilibria are always local minima with respect to director variations, but need not be minima of the full Landau–de Gennes energy.
More importantly, the intrinsic-basis analysis above shows that even a small departure from the Oseen–Frank manifold, $1-s\sim O(\varepsilon^{2})$, removes the bulk restoring force for director rotations and thereby permits destabilisation through coupling to magnitude fluctuations. This does not imply that every such deviation is unstable; rather, it implies that only a restricted class of Oseen–Frank solutions can persist as genuine equilibria of the full Landau–de Gennes theory, namely those for which this destabilising mechanism is exactly balanced by elastic and geometric effects.

### v6

@@ Abstract
Strong anchoring boundary conditions are conventionally modelled by imposing Dirichlet conditions on the order parameter in Landau–de Gennes theory, neglecting the finite surface energy of realistic anchoring. This work revisits the strong anchoring limit for nematic liquid crystals in confined two-dimensional domains. By explicitly retaining a Rapini–Papoular surface energy and adopting a scaling where the extrapolation length $l_{ex}$ is comparable to the coherence length $\xi$, we analyse both the small-domain ($\varepsilon=h/\xi\to 0$; $h$ is the domain size) and Oseen–Frank $(\varepsilon\to\infty$) asymptotic regimes. In the small-domain limit, the leading-order equilibrium solution is given by the average of the boundary data, which can vanish in symmetrically frustrated geometries, leading to isotropic melting. In the large-domain limit, matched asymptotic expansions reveal that surface anchoring introduces an $O(1/\varepsilon)$ correction to the director field, in contrast to the $O(1/\varepsilon^{2})$ correction predicted by Dirichlet conditions. The analysis captures the detailed structure of interior and boundary defects, showing that mixed (Robin-type) boundary conditions yield smoother defect cores and more physical predictions than rigid Dirichlet conditions. Numerical solutions for square and circular wells with tangential anchoring illustrate the differences between the two boundary condition treatments, particularly in defect morphology. These results demonstrate that a consistent treatment of anchoring energetics, together with stability considerations, is essential for accurate modelling of nematic equilibria in micro- and nano-scale confined geometries.
@@ 2 Remarks on the strong anchoring condition in three dimensions
To place these regimes in context, we note that in practice $l_{ex}\sim 10^{-8}-10^{-5}$[ravnik2009landau], with the strongest achievable anchoring corresponding to a few tens of nanometers. By contrast, the coherence length $\xi\sim 10^{-8}m$ [de1993physics] is typically only a few ångströms and rarely exceeds tens of nanometers even very close to the nematic–isotropic transition [^1] Strictly speaking, very close to the nematic–isotropic transition, fluctuation effects eventually invalidate the mean-field Landau–de Gennes description.. For example, using typical material parameters, $L=15pN$, $B=3.7\times 10^{5}Nm^{-2}$ and $C=2.4\times 10^{5}Nm^{-2}$ [rajamanickam2026colloidal], one finds $\xi=2.7\times 10^{-8}m$ at the transition point. Consequently, the regime $l_{ex}\ll\xi$ is seldom realised in practice. At this stage, it is worth noting that most studies in the literature impose Dirichlet conditions even within localized boundary defects corresponding to this infeasible regime.
@@ 3 Problem formulation in the reduced two-dimensional Landau–de Gennes framework
with $\boldsymbol{\nu}$ being the unit outward normal to $\partial\Omega$. The formulation for the associated stability analysis is provided in Appendix A.
@@ 4 Small-domain limit, $\varepsilon\to 0$
$q_{1}^{(0)}=\frac{1}{|\partial\Omega|}\oint_{\partial\Omega}q_{1,b}\,dl,\qquad q_{2}^{(0)}=\frac{1}{|\partial\Omega|}\oint_{\partial\Omega}q_{2,b}\,dl,\quad\Rightarrow\quad\mathbf{Q}^{(0)}=\frac{1}{|\partial\Omega|}\oint_{\partial\Omega}\mathbf{Q}_{b}\,dl.$ (18)
@@ 5 Oseen–Frank limit, $\varepsilon\to\infty$
Equilateral triangle [rajamanickam2026nematic] 44 three +1+1-vertex defects & one $\tfrac{1}{2}-$interior defect
@@
$\frac{\partial s}{\partial\nu}=-\varepsilon\gamma[s-\cos(2\varphi-2\varphi_{b})],\qquad 2s\frac{\partial\varphi}{\partial\nu}=-\varepsilon\gamma\sin(2\varphi-2\varphi_{b})\quad\text{on}\quad\partial\Omega.$ (33)
@@ 5.1 Outer solution
$s_{1}=0,\quad\varphi_{1}=\frac{1}{4\gamma}\frac{\partial\varphi_{0}}{\partial\nu}\quad\text{on}\quad\partial\Omega.$ (38)
@@
$s_{2}=-\frac{1}{2\gamma^{2}}\left(\frac{\partial\varphi_{0}}{\partial\nu}\right)^{2},\quad\varphi_{2}=\frac{1}{4\gamma}\frac{\partial\varphi_{1}}{\partial\nu}\quad\text{on}\quad\partial\Omega.$ (40)
@@
The first-order correction exists only because $\gamma\sim O(1)$. In studies where surface energy is neglected $(\gamma\to\infty)$, $\varphi_{1}=0$ and the first correction vanishes. Properly accounting for surface energy introduces an $O(1/\varepsilon)$ correction to the director field, in contrast to the $O(1/\varepsilon^{2})$ prediction from purely Dirichlet boundary conditions [nguyen2013refined, di2020landau]. The bulk scalar order parameter, however, satisfies
$s=1-\frac{2}{\varepsilon^{2}}|\nabla\varphi_{0}|^{2}-\frac{4}{\varepsilon^{3}}\nabla\varphi_{0}\cdot\nabla\varphi_{1}-\cdots.$ (41)
The director field $\mathbf{n}=(\cos\varphi,\sin\varphi)^{T}$ admits the expansion
@@ 5.3 Structure of a boundary defect
$\frac{2S_{0}}{\gamma\eta}\frac{\partial\Phi_{0}}{\partial\theta}=\sin(2\Phi_{0}-2\varphi_{b})\quad\text{on}\quad\theta=\pm\alpha,$ (67)